Factor the following
x^2+-9+1x
bx^2+-7+-5
x^2+-36
3x^2+11x+4
x^2+8x+7
5x^2+x-18
9x^2+bx-8
26x+x^2-18

Answer by stanbon(75887)

Factor the following
x^2+-9+1x
= x^2+x-9
x = [-1 +- sqrt(1 - 4*-9)]/2
x = [-1 +- sqrt(37)]/2
x = (-1/2)+ (1/2(sqrt37) or x = (-1/2)-(1/2)sqrt(37)
(x-(-1/2)+ (1/2(sqrt37))(x(-1/2)- (1/2(sqrt37))
------------------------------------
bx^2+-7+-5
Same pattern as above using the quadratic formula.
-----------------
x^2+-36
=(x-6)(x+6)
-----------------
3x^2+11x+4
Use the quadratic formula.
---------------------------
x^2+8x+7
=(x+7)(x+1)
---------------
5x^2+x-18
= 5x^2+10x-9x-18
= 5x(x+2)-9(x+2)
= (x+2)(5x-9)
------------------
9x^2+bx-8
Use the quadratic formula.
-------------------
26x+x^2-18
x^2+26x-18
Use the quadratic formula
=============================
Cheers,
Stan H.